If y₁ = 2x+2 and y₂= -x²/2 are the solutions of the equation y= xy´+( y´)²/2 ,
Then are the constant multiples c₁y₁ and c₂y₂ , where c₁,c₂ are arbitary, also the solutions of the given DE ? Is the sum y₁+y₂ a solution ? Justify your answer .
1
Expert's answer
2019-08-20T09:43:37-0400

1)Let's check, if "c_1y_1" is a solution:

"c_1y_1=2c_1(x+1)." .

Let's check, if "2c_1(x+1)" is equal to "x*(2c_1(x+1))'+\cfrac{((2c_1(x+1))')^2}{2}" :

"x*(2c_1(x+1))'+\cfrac{((2c_1(x+1))')^2}{2} = 2c_1x+2c_1^2" and it's not equal to "2c_1(x+1)" for arbitary "c_1" so it's not a solution.

2)Let's check if "c_2y_2" is a solution:

"c_2y_2=\cfrac{-c_2x^2}{2}" .

Let's check, if "\cfrac{-c_2x^2}{2}" is equal to "x*(\cfrac{-c_2x^2}{2})'+\cfrac{((\cfrac{-c_2x^2}{2})')^2}{2}:"

"x*(\cfrac{-c_2x^2}{2})'+\cfrac{((\cfrac{-c_2x^2}{2})')^2}{2}=-c_2x^2+\cfrac{c_2^2x^2}{2}" and it's not equal to "\cfrac{-c_2x^2}{2}" for arbitary "c_2" so it's not a solution.

3)Let's check, if "y_1+y_2" is a solution:

"(y_1+y_2)'=(\cfrac{-x^2}{2}+2x+2)'=2-x" .

So "xy´+\cfrac{( y´)²}{2}=2x-x^2+\cfrac{1}{2}" and it's not equal to "y_1+y_2" .

Answer: none of the three expressions is a solution.